OpenStudy (elisaneedshelp):

Solve the following quadratic for x by any method - Completing the Square, Factoring or Quadratic Formula. x2 + 10x + 24 = 0

OpenStudy (elisaneedshelp):

could you please help me with this one too if you can @jigglypuff314 :)

jigglypuff314 (jigglypuff314):

what two numbers multiply to get 24 and also add to get 10 ? :)

OpenStudy (elisaneedshelp):

6 and 4, which method are you using? lol

jigglypuff314 (jigglypuff314):

factoring :P

OpenStudy (elisaneedshelp):

I have to explain this to my teacher over the phone so I just say that I used factoring Awesomeee, so those are both solutions? :P

OpenStudy (elisaneedshelp):

Can we try another method because I don't get how to show my work and they want me to show the work :(

OpenStudy (elisaneedshelp):

I hope you're still there lol, if you hate the quadratic formula as much as me you might be annoyed but can you help me do it that way because showing my work for factoring blew my mind and it confuses me

jigglypuff314 (jigglypuff314):



jigglypuff314 (jigglypuff314):

sorry about the wait - I got side tracked by a trouble maker .-. anyways... x^2 + 10x + 24 = 0 a = 1 (the coefficient in front of x^2) b = 10 c = 24

OpenStudy (elisaneedshelp):

Lol it's fine I was just thinking if you had an emergency irl or something and had to leave me lol or bathroom break

jigglypuff314 (jigglypuff314):

so we can just plug in those a b c into the quadratic formula

OpenStudy (kayders1997):

why not factor? its easier

jigglypuff314 (jigglypuff314):

factoring is more difficult to explain conceptually ^_^

OpenStudy (kayders1997):

oh I see

OpenStudy (kayders1997):

you could think of it as what times what equal x^2 and what times what multiply to 24 and adds to 10

OpenStudy (kayders1997):

you could think of it as what times what equal x^2 and what times what multiply to 24 and adds to 10

OpenStudy (elisaneedshelp):